(b) Global Insurance Company has four salespeople working in Hill town. The number of policies sold during the last month is given in matrix A, as 


o1 Malley Coponi Steomberg Deutsch

8 7 6 8 Automobile

6 9 11 5 Life

A= 4 3 2 0 Health

0 2 1 3 Homeowners


i) Let S= [1 1 1 1]. Find SA and interpret its elements.

ii) Find 𝐴𝑆௧ and interpret its elements. 


1
Expert's answer
2021-06-16T15:55:12-0400

"A=\begin{bmatrix}8&7&6&8\\6&9&11&5\\4&3&2&0\\0&2&1&3\end{bmatrix}"


"(1) SA=\begin{bmatrix} 1&1&1&1\end{bmatrix} \begin{bmatrix}8&7&6&8\\6&9&11&5\\4&3&2&0\\0&2&1&3\end{bmatrix} 



 =\begin{bmatrix} 18&21&20&16\end{bmatrix}"


       Elements are 18,21,20 and 16.In this case we get a horizontal matrix of order 1x4.


"(2)

 AS^T=\begin{bmatrix}8&7&6&8\\6&9&11&5\\4&3&2&0\\0&2&1&3\end{bmatrix} \begin{bmatrix} 1\\1\\1\\1\end{bmatrix}



 =\begin{bmatrix} 18\\21\\20\\16\end{bmatrix}"


    Its elements are same as the above AS matrix.In this we get a vertical matrix of order 4 x 1.